# Partial Resolution of the Erdős-Straus, Sierpiński, and Generalized Erdős-Straus Conjectures Using New Analytical Formulas

### v1

@@ 25 R = 1/x + 1/y + 1/z # Right-hand side of the equation
30
31 # Example usage:
32 a = 4 #(for the classic Erdos_Strauss Conjecture)
33 n_min = 2
34 n_max = 100
35 generalized_erdos_straus_conjecture(a, n_min, n_max)
@@
1 def second_formula_erdos_straus(a, n_min, n_max):
@@
4 This function prints results directly based on the existence of a perfect square.
5 """
6 epsilon = 1e-9 # Tolerance for floating-point comparison
7 for n in range(n_min, n_max + 1):
8 L = a / n # Left-hand side of the equation
9 for x in range(n // a + 1, 10 * n): # Searching for x > n/a
10 t_min = max(1, (2 * n * x) // ((a * x - n) ** 2)) # Lower bound for t
11 for t in range(t_min, t_min + 100): # Iterate over possible t values in a reasonable range
12 q_square = t**2 * (a*x - n)**2 - 2 * n * t*x
13 if q_square >= 0:
14 q = int(q_square**0.5)
15 if q * q == q_square: # Ensure q is a perfect square
16 y = t * (a*x - n) - q
17 z = t * (a*x - n) + q
@@ 8 L = a / n # Left-hand side of the equation
19 R = 1/x + 1/y + 1/z
20 is_true = abs(L - R) < epsilon
21 print(f"For n = {n}, x = {x}, y = {y}, z = {z}, t = {t}: L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
@@ 19 R = 1/x + 1/y + 1/z
23
24 # Example usage:
25 a = 4#(a=4 for the classic erdos_strauss conjecture,we can change de value of a>=4 depending on the conjecture we want to test)
26 n_min = 2
27 n_max = 150
28 #for a fixed value of a , we can vary the interval[n_min,n_max] to extend n.
29 print("Testing Second Formula:")
30 second_formula_erdos_straus(a, n_min, n_max)
Appendix : Python Code for Testing the classic Erdős-Straus Conjecture for the Values Not Proven by Mordell (Second Formula)
1 def second_formula_erdos_straus_mordell(a, n_min, n_max):
2 """
@@
4 for values of n where Mordell’s method was inconclusive.
5 """
6 epsilon = 1e-9 # Tolerance for floating-point comparison
7 mordell_exceptions_mod_840 = {1, 121, 169, 289, 361, 529, 841, 961, 1009, 1129, 1369, 1681, 1849, 1969} # Corrected list of exceptional cases
8 captured_cases = 0 # Counter for validated cases
9 total_cases = 0 # Total exceptional cases in range
10
11 print("Mordell’s exceptional cases in the given range:")
12 for n in range(n_min, n_max + 1):
13 if n % 840 in mordell_exceptions_mod_840:
14 total_cases += 1
15 print(n, end=", ")
@@
17
18 print("Testing Second Formula for Mordell’s Exceptional Cases:")
19 for n in range(n_min, n_max + 1):
20 if n % 840 not in mordell_exceptions_mod_840:
21 continue # Ignore values where Mordell already proved the conjecture
22
23 L = a / n # Left-hand side of the equation
24 solution_found = False
25
26 for x in range(n // a + 1, 10 * n): # Searching for x > n/a
27 t_min = max(1, (2 * n * x) // ((a * x - n) ** 2)) # Lower bound for t
28 for t in range(t_min, t_min + 100): # Iterate over possible t values in a reasonable range
29 q_square = t**2 * (a*x - n)**2 - 2 * n * t*x
30 if q_square >= 0:
31 q = int(q_square**0.5)
32 if q * q == q_square: # Ensure q is a perfect square
33 y = t * (a*x - n) - q
34 z = t * (a*x - n) + q
35 if y > 0 and z > 0:
36 R = 1/x + 1/y + 1/z
37 is_true = abs(L - R) < epsilon
38 captured_cases += 1
39 print(f"For n = {n} (Mordell case), x = {x}, y = {y}, z = {z}, t = {t}: "
40 f"L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
41 solution_found = True
42 break
43 if solution_found:
44 break
45
46 if total_cases > 0:
47 percentage_captured = (captured_cases / total_cases) * 100
@@ 36 R = 1/x + 1/y + 1/z
50 print("No exceptional cases found in the given range.")
51
52 # Example usage:
@@
54 n_min = 2
55 n_max = 2000 # Increase to test larger cases
56
57 print("Testing Second Formula for Mordell’s Exceptional Cases:")
58 second_formula_erdos_straus_mordell(a, n_min, n_max)
@@
Appendix: python code for percentages of n captures and recalcitrant cases
1
2
3 def second_formula_erdos_straus(a, n_min, n_max):
4 """
[…]
6 Stops searching for a given n as soon as one solution is found.
7 Computes the percentage of n captured and lists the uncaptured n values.
8 """
9 epsilon = 1e-9 # Tolerance for floating-point comparison
10 captured_n = 0 # Counter for the number of n with at least one solution
11 total_n = n_max - n_min + 1 # Total n values in the range
12 uncaptured_ns = [] # List to store n values that were not captured
13
14 for n in range(n_min, n_max + 1):
15 L = a / n # Left-hand side of the equation
16 found = False # Flag to check if a solution is found for this n
17
18 for x in range(n // a + 1, 100 * n): # Searching for x > n/a
19 t_min = max(1, (2 * n * x) // ((a * x - n) ** 2)) # Lower bound for t
20
21 for t in range(t_min, t_min + 100): # Iterate over possible t values
22 q_square = t**2 * (a * x - n)**2 - 2 * n * t * x
23 if q_square >= 0:
24 q = int(q_square**0.5)
25 if q * q == q_square: # Ensure q is a perfect square
26 y = t * (a * x - n) - q
27 z = t * (a * x - n) + q
28 if y > 0 and z > 0:
29 R = 1/x + 1/y + 1/z
30 is_true = abs(L - R) < epsilon
31 print(f"For n = {n}, x = {x}, y = {y}, z = {z}, t = {t}: L = {L:.4f}, R = {R:.4f}, Conjecture holds: {is_true}")
32
33 captured_n += 1 # Increment the counter since we found a solution
34 found = True
35 break # Stop searching for more solutions for this n
36
37 if found:
38 break # Stop searching for this n as we already found a solution
39
40 if not found:
41 uncaptured_ns.append(n) # Store n if no solution was found
42
43 # Compute and print the percentage of captured n
44 percentage_captured = (captured_n / total_n) * 100
[…]
46
47 # Print uncaptured n values
48 if uncaptured_ns:
[…]
50 print(uncaptured_ns)
51 else:
[…]
53
54 # Example usage:
[…]
56 n_min = 2
57 n_max = 100
58 print("Testing Second Formula:")
59 second_formula_erdos_straus(a, n_min, n_max)
Appendix: python code for finding perfect square manually (Case $n=577$)
1
2 import math
3
4 def find_perfect_squares(start=18592, end=20000):
5 results = []
6 for t in range(start, end + 1): # Iterate over t in a reasonable interval
7 delta = t * (9 * t - 167330) # Correction here with 9 * t
8 if delta > 0: # Ensure delta is positive
9 sqrt_delta = math.isqrt(delta) # Integer square root of delta
10 if sqrt_delta * sqrt_delta == delta: # Check if it’s a perfect square
11 results.append((t, delta, sqrt_delta))
12
13 return results
14
15 # Execute the code
16 start_value = 25000
17 end_value = 40000 # Define a reasonable upper bound here
18 perfect_squares = find_perfect_squares(start=start_value, end=end_value)
19
20 # Display results
21 for t, delta, square_root in perfect_squares:
22 print(f"t = {t}, delta = {delta}, racine_delta = {square_root}")
23
24 if not perfect_squares:
25 print(f"No values of t between {start_value} and {end_value} produce a perfect square.")

### v2

@@ 27 R = 1/x + 1/y + 1/z # Right-hand side of the equation
32
33# Example usage:
34a = 4 #(for the classic Erdos_Strauss Conjecture)
35n_min = 2
36n_max = 100
37generalized_erdos_straus_conjecture(a, n_min, n_max)
@@
1import math
2import time
3from fractions import Fraction
4
5print("VERIFICATION OF OUR CONJECTURE -- CASE a = 4")
6print("="*70)
7
8def verify_our_conjecture(n_min=2, n_max=500,
9 max_x_factor=300,
10 max_t_range=500,
11 max_q=10):
12
13 captured = 0
14 failed_n_list = []
15
16 start_time = time.time()
17 a = 4# You can change de value of a to test another conjecture
18
19 for n in range(n_min, n_max + 1):
20 found = False
21
22 x_start = n // a + 1
23 x_limit = max_x_factor * n
24
25 for x in range(x_start, x_limit + 1):
26 d = a * x - n
27 if d == 0:
28 continue
29
30 denom = d * d
31 t_min = max(1, (2 * n * x) // denom)
32
33 # =========================
34 # PHASE 1 : integer t
35 # =========================
36 for t in range(t_min, t_min + max_t_range + 1):
37 F = t*t*d*d - 2*t*x*n
38 if F < 0:
39 continue
40
41 q = math.isqrt(F)
42 if q*q == F:
43 y = t*d - q
44 z = t*d + q
45 if y > 0 and z > 0 and y == int(y) and z == int(z):
46 if a*x*y*z == n*(y*z + x*z + x*y):
47 captured += 1
48 found = True
49 print(f"n={n:4d} | x={x}, t={t} [INTEGER], y={y}, z={z}, F={F}, q={q}")
50 break
51 if found:
52 break
53
54 # =========================
55 # PHASE 2 : rational t
56 # =========================
57 if not found:
58 for x in range(x_start, x_limit + 1):
59 d = a * x - n
60 if d == 0:
61 continue
62
63 for q_den in range(2, max_q+1):
64 for p in range(1, max_q*30):
65 t = Fraction(p, q_den)
66 F = t*t*d*d - 2*t*x*n
67 if F < 0:
68 continue
69
70 F_num = F.numerator
71 F_den = F.denominator
72 if math.isqrt(F_num)**2 == F_num and math.isqrt(F_den)**2 == F_den:
73 qrat = Fraction(math.isqrt(F_num), math.isqrt(F_den))
74 y = t*d - qrat
75 z = t*d + qrat
76 # === require y and z to be integers ===
77 if y.denominator == 1 and z.denominator == 1 and y > 0 and z > 0:
78 if a*x*int(y)*int(z) == n*(int(y)*int(z) + x*int(z) + x*int(y)):
79 captured += 1
80 found = True
81 print(f"n={n:4d} | x={x}, t={t} [RATIONAL], y={y}, z={z}, F={F}, q={qrat}")
82 break
83 if found:
84 break
85 if found:
86 break
87
88 if not found:
89 print(f"n={n:4d} : X FAILED")
90
91 failed_n_list.append(n)
92
93 elapsed_time = time.time() - start_time
94 total = n_max - n_min + 1
95 success_rate = 100 * captured / total
96
[…]
98 print("FINAL REPORT a = 4")
99 print("="*70)
100 print(f"Tested interval : n in [{n_min}, {n_max}]")
101
102 print(f"Success : {captured}/{total} ({success_rate:.2f}%)")
103 print(f"Time : {elapsed_time:.2f}s")
104
105 if failed_n_list:
106 print("Detected failures :", failed_n_list)
107 else:
108 print(" COMPLETE SUCCESS")
109
110 return
111
112
113
114# DIRECT EXECUTION
115
116
117verify_our_conjecture(n_min=2, n_max=500)#You can vary the check interval for a fixed a
Appendix: Python Code for Testing the Classic Erdos-Straus Conjecture for the Values Not Proven by Mordell (Second Formula)
1
2import math
3import time
4
5print("VERIFICATION OF ERDOS-STRAUS CONJECTURE - a = 4 (Our Method)")
6
7print("Testing n in quadratic residues modulo 840")
8print("=" * 70)
9
10# Parameters
11a = 4
12MOD = 840
13RESIDUES = {(r * r) % MOD for r in [1, 11, 13, 17, 19, 23]}
14
15def generate_admissible_n(n_min, n_max):
16 """Generate list of admissible n (quadratic residues mod 840) in [n_min, n_max]"""
17 admissible_list = []
18 for n in range(n_min, n_max + 1):
19 if (n % MOD) in RESIDUES:
20 admissible_list.append(n)
21 return admissible_list
22
23def verify_with_our_method_for_list(n_list, max_x_factor=300, max_t_range=500):
@@ 66 F = t*t*d*d - 2*t*x*n
25 Verify the conjecture using our method for a given list of n.
26 Returns: solutions dict, failures list, elapsed time
27 """
28
29 captured = 0
30 failed_n_list = []
31 solutions = {}
32
33 start_time = time.time()
34 total_n = len(n_list)
35
@@ 26 Returns: solutions dict, failures list, elapsed time
37 print(f"Parameters: max_x_factor={max_x_factor}, max_t_range={max_t_range}")
38 print("-" * 70)
39
40 for idx, n in enumerate(n_list, 1):
41 found = False
42
43 # Search for x
44 x_start = n // a + 1
45 x_limit = max_x_factor * n
46
47 for x in range(x_start, x_limit + 1):
48 if found:
49 break
50
51 d = a * x - n
52 if d == 0:
53 continue
54
55 # Compute t_min to ensure F >= 0
56 denom = d * d
57 t_min = max(1, (2 * n * x) // denom)
58
59 # Search for t
60 for t in range(t_min, t_min + max_t_range + 1):
61 F = t * t * d * d - 2 * t * x * n
62
63 if F < 0:
64 continue
65
66 q = math.isqrt(F)
67 if q * q == F:
68 y = t * d - q
69 z = t * d + q
70
@@ 61 F = t * t * d * d - 2 * t * x * n
72 left = a * x * y * z
73 right = n * (y*z + x*z + x*y)
74
75 if left == right:
76 solutions[n] = (x, t, y, z)
77 captured += 1
78 found = True
79
80 # Print solution
81 print(f"n={n:4d} : 0k x={x:5d}, t={t:5d}, y={y:6d}, z={z:6d}")
@@
83
84 if not found:
85 failed_n_list.append(n)
86 print(f"n={n:4d} : FAILED")
87
88
89 elapsed_time = time.time() - start_time
90
91 return solutions, failed_n_list, elapsed_time
92
93def intensive_search(failed_list, max_x_factor=1000, max_t_range=2000):
94 """Intensive search for failed n"""
95
@@
97 print(f"INTENSIVE SEARCH FOR {len(failed_list)} FAILED N")
98 print("="*70)
99
100 a = 4
101 new_solutions = {}
102
103 for idx, n in enumerate(failed_list, 1):
104 print(f"Failed {idx}/{len(failed_list)} - n={n} : ", end="", flush=True)
105
106 found = False
107 x_start = n // a + 1
108 x_limit = max_x_factor * n
109
110 for x in range(x_start, min(x_limit, 50000) + 1):
111 if found:
112 break
113
114 d = a * x - n
115 if d <= 0:
116 continue
117
118 t_min = max(1, (2 * n * x) // (d * d))
119 t_max = t_min + max_t_range
120
121 for t in range(t_min, min(t_max, t_min + 5000) + 1):
122 F = t * t * d * d - 2 * t * x * n
123
124 if F < 0:
125 continue
126
127 q = math.isqrt(F)
128 if q * q == F:
129 y = t * d - q
130 z = t * d + q
131
132 if y > 0 and z > 0:
133 left = a * x * y * z
134 right = n * (y*z + x*z + x*y)
135
136 if left == right:
137 new_solutions[n] = (x, t, y, z)
138 found = True
139 print(f"ok (x={x}, t={t}, y={y}, z={z})")
140 break
141
142 if not found:
143 print("FAILED")
144
145
146 return new_solutions
147
148# ========================================
149# MAIN EXECUTION
150# ========================================
151if __name__ == "__main__":
152 # MODIFY THESE VALUES TO TEST DIFFERENT INTERVALS
153 n_min = 5000 # YOU CAN CHANGE THIS
154 n_max = 10000 # YOU CAN CHANGE THIS
155
156 # Step 1: Generate admissible n in the interval
@@
158 print("-" * 70)
159
160 admissible_list = generate_admissible_n(n_min, n_max)
161
162 if len(admissible_list) == 0:
163 print("No admissible n in this interval!")
164 exit()
165
166 print(f"Total admissible n: {len(admissible_list)}")
[…]
168 # Print admissible n in groups of 20 for readability
169 for i in range(0, len(admissible_list), 20):
170 print(f" {admissible_list[i:i+20]}")
171
172 # Step 2: Verify using our method
173 solutions, failed, elapsed = verify_with_our_method_for_list(
174 n_list=admissible_list,
175 max_x_factor=300,
176 max_t_range=500
177 )
178
179 # Step 3: Final report
[…]
181 print("FINAL REPORT")
182 print("="*70)
183
184 success_rate = 100 * len(solutions) / len(admissible_list)
185
186 print(f"Interval tested : [{n_min}, {n_max}]")
187 print(f"Admissible n tested : {len(admissible_list)}")
188 print(f"Execution time : {elapsed:.2f} seconds")
189 print(f"Solutions found : {len(solutions)}/{len(admissible_list)} ({success_rate:.2f}%)")
190 print(f"Failures : {len(failed)}")
191
192 if failed:
[…]
194 for i in range(0, len(failed), 20):
195 print(f" {failed[i:i+20]}")
196
197 # Step 4: Intensive search for failures
198 new_solutions = intensive_search(
199 failed_list=failed,
200 max_x_factor=1000,
201 max_t_range=2000
202 )
203
204 if new_solutions:
[…]
206 print(f"INTENSIVE SEARCH RESULTS")
207 print("="*70)
208
209 print(f"New solutions found: {len(new_solutions)}")
210 for n in sorted(new_solutions.keys()):
211 x, t, y, z = new_solutions[n]
212 print(f" n={n:4d} : x={x:6d}, t={t:6d}, y={y:8d}, z={z:8d}")
213
214 # Update statistics
215 solutions.update(new_solutions)
216 remaining_failures = [n for n in failed if n not in new_solutions]
217 new_total = len(solutions)
218 new_rate = 100 * new_total / len(admissible_list)
219
[…]
221 print(f" Total solutions : {new_total}/{len(admissible_list)} ({new_rate:.2f}%)")
222 print(f" Remaining failures : {len(remaining_failures)}")
223
224 if remaining_failures:
225 print($\nRemaining f$ailures ({len(remaining_failures)} numbers):")
226 for i in range(0, len(remaining_failures), 20):
227 print(f" {remaining_failures[i:i+20]}")
228 else:
@@
232
[…]
234 print("END OF PROGRAM")
235 print("="*70)
236
237 # Return the admissible list as well (useful for further analysis)
238 admissible_list_result = admissible_list